OpenStudy (firejay5):

A rectangle has a perimeter of 30 centimeters. Which of the following could NOT be the area for this rectangle? A. 26 square centimeters B. 50 square centimeters C. 54 square centimeters D. 60 square centimeters

OpenStudy (firejay5):

@agent0smith

OpenStudy (triciaal):

perimeter of rectangle = 2(l + w) = 30 so l + w = 15 length = l and width = w area = l*w if the l and w were approximately the same 15/2 get 7 and 8 9 and 6 area = 54 10 and 5 area 50 12 and 3 area the area could not be 26

OpenStudy (agent0smith):

@triciaal 2 and 13 gives an area of 26

OpenStudy (agent0smith):

60 doesn't seem possible

OpenStudy (firejay5):

why you say that

OpenStudy (agent0smith):

The greatest area will be when it's a square, so side length 7.5, which gives a max possible area of 56.25 cm^2, so anything below that is valid I think

OpenStudy (triciaal):

@agent smith thank you I forgot the logic and then when substituting i did opposite \ @Firejay5 sorry correct answer as given by @agent0smith

OpenStudy (firejay5):

is it A or D

OpenStudy (triciaal):

d the maximum area is 56 when the l = w so it cannot be 60